# Synchronization of two coupled phase oscillators in the time-delayed second-order Kuramoto model

### v1

@@ III.2 Non-Identical oscillators
Our study also illustrated that the presence of time-delayed interaction in the second-order Kuramoto model can generate non-phase-locked solutions, which can lead to limit-cycle or chaotic behaviors based on the inertia coefficient and the amount of time delay. Our findings suggested that the observed behavior is highly sensitive to the value of the time delay. Even a slight change in the time delay can cause a shift from periodic to chaotic behavior.
For systems with significant inertia and time delay, behaviors resembling epileptic seizures can be observed in the system’s dynamics. These behaviors occur when the oscillators are nearly fully synchronized for a period of time before eventually going out of phase. A previous study by Gerster et al. Gerster et al. (2020) reported the spontaneous occurrence of synchronization phenomena that closely resemble the observed behavior during epileptic seizures. This occurrence is attributed to the complex interaction between dynamics and network topology. The impact of time delay is significant in the context of epilepsy, as seizures are often observed as a symptom in demyelinating diseases, such as multiple sclerosis de Curtis, Garbelli, and Uva (2021). These observations suggest a potential connection between the dynamics of time-delayed coupled oscillators and the pathological synchronization observed in epileptic seizures and demyelinating diseases. However, further research is needed to explore this connection.
Our findings show that when there is a difference in frequency, it expands the stability range of fixed point solutions to longer time delay values. Additionally, in the case of non-fixed point solutions, the increase in frequency disparity can cause a shift from chaotic to regular dynamics.
Overall, the study of coupled time-delayed oscillators provides a valuable framework for understanding the complex behaviors observed in many natural and engineered systems. Therefore, we hope that our findings will contribute to a deeper understanding of the underlying mechanisms that govern their behavior.

### v2

@@ III.2 Case $\omega_{1}\neq\omega_{2}$
Our study also illustrated that the presence of time-delayed interaction in the second-order Kuramoto model can generate non-phase-locked solutions, which can lead to limit-cycle or chaotic behaviors based on the inertia coefficient and the amount of time delay. Our findings suggested that the observed behavior is highly sensitive to the value of the time delay. Even a slight change in the time delay can cause a shift from periodic to chaotic behavior. Our findings also show that the difference in interinsic frequencies expands the stability range of fixed point solutions to longer time delay values. Additionally, in the case of non-fixed point solutions, the increase in frequency disparity can cause a shift from chaotic to regular dynamics.
For systems with significant inertia and time delay, behaviors resembling epileptic seizures can be observed in the system’s dynamics. These behaviors occur when the oscillators are nearly fully synchronized for a period of time before eventually going out of phase. A previous study by Gerster et al. Gerster et al. (2020) reported the spontaneous occurrence of synchronization phenomena that closely resemble the observed behavior during epileptic seizures. This occurrence is attributed to the complex interaction between dynamics and network topology. The impact of time delay is significant in the context of epilepsy, as seizures are often observed as a symptom in demyelinating diseases, such as multiple sclerosis de Curtis, Garbelli, and Uva (2021). Indeed, the transmission of information in neuronal systems is influenced by several factors, one of which is the presence of myelin. Myelin is a fatty substance that encircles axons, providing an insulating layer that boosts the speed and efficiency of signal propagation. If myelin is lost or damaged, such as in certain neurological disorders or injuries, the transmission of signals can slow down significantly, leading to a longer time delay between the interactions of neurons. These observations suggest a potential connection between the dynamics of time-delayed coupled massive oscillators and the pathological synchronization observed in epileptic seizures and demyelinating diseases.
It has been shown that the Josephson junctions can be used to model the dynamics of a single neuron Mishra et al. (2021). A simplified model of a neuron with an axon and dendrites can be represented as a two-dimensional phase oscillator with inertia. This model is capable of explaining the intrinsic transient behavior commonly observed in experimental settings of the networks of dendritic neurons Dolan, Majtanik, and Tass (2005); Majtanik, Dolan, and Tass (2006); Sakyte and Ragulskis (2011). Our results also indicate that the second-order Kuramoto model may be more suitable for describing the behavior of neurons that the original Kuramoto model cannot capture. Further research is necessary to explore this connection.
Overall, the study of coupled time-delayed massive oscillators provides a valuable framework for understanding the complex behaviors observed in many natural and engineered systems. Therefore, we hope that our findings will contribute to a deeper understanding of the underlying mechanisms that govern their behavior.